OpenStudy (anonymous):

**MEDAL** I'm making flashcards to help me with my ASVAB. Help me with some formulas? Davis donates 4/13 of his paycheck to his favorite charity. If he donates $26.80, what is the amount of his paycheck? A. $8.25 B. $82.50 C. $87.10 <-- D. $348.40 Formula?

OpenStudy (anonymous):

4/13 times (paycheck) = $26.80 (paycheck) = 13/4 times $26.80

OpenStudy (anonymous):

Thanks, @heavenfun13

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

26.8/4 is 1/13 of this whole paycheck so simply multiply it by another 13 to get the total amount. 26.8/4= 6.7*13=87.1